Given that log b (xy^3)=m and log b (x^3y^2)=p, find log b ([square root]xy) in terms of m and p.
Answer by AnlytcPhil(1116)

Given that log b (xy^3)=m and log b (x^3y^2)=p, find log b ([square root]xy)
in terms of m and p.
 
The three basic principles of logarithms are

(1) logB(MN) = logM + logN

(2) logB(M/N) = logBM + logBN

(3) logBMN = N·logBM

We'll use these three rules on the equations for m and p:

           logb(xy3) = m

Use (1) on left side:

      logbx + logby3 = m  

Use (3) on the second term:

     logbx + 3·logby = m

-------------------------------

Next:

          logb(x3y2) = p 

Use (1)

     logbx3 + logby2 = p       

Use (3) on each of the terms on the left:

   3·logbx + 2·logby = p

-------------------------------------------

     logbx + 3·logby = m
   3·logbx + 2·logby = p

This is a system of two equations and two unknowns, so we 
solve it by Cramer's rule for logbx and logby

            |m 3|
|p 2| 2m-3p 2m-3p -3p+2m 3p-2m logbx = ------- = ------- = ------- = --------- = ------- |1 3| 2-9 -7 -7 7 |3 2| |1 m| |3 p| p-3m p-3m -3m+p 3m-p logby = ------- = ------ = ------ = ------- = ------ |1 3| 2-9 -7 -7 7 |3 2| Now we use the rules on the expression: __ logbÖxy The square root is the same as the 1/2 power, so we write: logb(xy)1/2 Use (3) (1/2)·logb(xy) Use (1) on the log (1/2)·(logbx + logby) Now replace logbx by (3p-2m)/7 and logy by (3m-p)/7 (1/2)·[(3p-2m)/7 + (3m-p)/7] (1/2)(1/7)·[(3p-2m) + (3m-p)] (1/14)·[3p-2m+3m-p] (1/14)·(2p+m) (2p+m)/14 Edwin J